OpenStudy (anonymous):

suppose you invest 1050 at an annual interest rate of 5.5% compounded continously how much will you have in the account after 5 years round the solution to the nearest dollar

OpenStudy (anonymous):

hmm let me think

OpenStudy (anonymous):

A=P(1+r/n)^tn P = principal amount (the initial amount you borrow or deposit - $1050) r = annual rate of interest (in decimal - .055) t = number of years the amount is deposited or borrowed for. n = number of times the interest is compounded per year

OpenStudy (anonymous):

P=I/RT

OpenStudy (anonymous):

does this help?